OpenStudy (anonymous):

n/5 greater then sign with the line under it 11

OpenStudy (zarkon):

\[\frac{n}{5}\ge 11\] \frac{n}{5}\ge 11

OpenStudy (anonymous):

i dont know if its n over 5 or not

OpenStudy (ksaimouli):

yup its is over 5

OpenStudy (anonymous):

so the answer would be 11?

OpenStudy (ksaimouli):

nope 11*5 =55